OpenStudy (anonymous):

Easy Question PLEASE HELP! 6. Simplify the expression: (4x^2 - 2x - 1) - (-3x^3 + 2)

OpenStudy (anonymous):

@pooja195 Please help.

OpenStudy (anonymous):

@helpmewithmathplzzzz

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

This is not the answer to your question just an example that might help you I do not want you to be dependent and please do not call me mean 3 + 2[4x – (4 + 3x)] = –1 3 + 2[4x – 1(4 + 3x)] = –1 3 + 2[4x – 1(4) – 1(3x)] = –1 3 + 2[4x – 4 – 3x] = –1 3 + 2[1x – 4] = –1 3 + 2[1x] + 2[–4] = –1 3 + 2x – 8 = –1 2x + 3 – 8 = –1 2x – 5 = –1 2x – 5 + 5 = –1 + 5 2x = 4 x = 2